# A Survey on Collaborating Small and Large Language Models for Performance, Cost-effectiveness, Cloud-edge Privacy, and Trustworthiness

### v1

@@ 3.1 Collaboration During Pre-Training Stage
Pre-training is the most computationally intensive phase of the LLM lifecycle, motivating cost-effective strategies. SLM–LLM collaboration improves data efficiency and optimizes training via three approaches: (1) data curation, where LLMs refine SLM data (Fig.4(a)) or SLMs assess LLM data quality (Fig.4(b)); (2) co-curriculum learning (Fig.5(a)), which progressively expands model capacity; and (3) pre-training distillation (Fig.5(b)), transferring knowledge from teacher to student models. (1) Data Curation
(a) SLM selects/Reshape Data and feeds into LLM
@@
High-quality data is essential for effective pre-training, yet real-world corpora are often noisy, inflating computational cost and degrading performance. SLM–LLM collaboration addresses this by filtering and refining datasets to improve data efficiency. For example, SALT Rawat et al. (2024) uses compact models to provide soft labels and select informative samples, accelerating LLM pre-training while reducing FLOPs and energy. Conversely, Ankner et al. (2024) employ perplexity-based pruning with SLM scorers to remove low-yield data, reducing optimization steps and enhancing generalization.
@@
Figure 5: (a) Co-curriculum learning and (b) Knowledge distillation.
LiGO Wang et al. (2023) maps small-model parameters to larger ones via a linear growth operator, cutting pre-training compute by up to 50%. Progressive training Yano et al. (2025) incrementally scales width and depth, saving $\sim$25% compute. HyperCloning Samragh et al. (2024) clones linear layers for stable warm starts and substantial GPU-hour reductions.
(3) Pre-Training Distillation Pre-training distillation (PD) extends classic task-agnostic methods such as DistilBERT, TinyBERT, and MiniLM Sanh et al. (2019); Jiao et al. (2020); Wang et al. (2020), now scaled to LLMs. Recent studies show that larger students can sometimes outperform teachers Peng et al. (2025) and disentangle PD effects from data confounds Goyal et al. (2025). MINIPLM Gu et al. (2025) refines data with teacher inference before student training, enhancing efficiency and diversity.
@@ 3.2 Collaboration During Tuning Stage
LLM tuning is highly resource-intensive, demanding substantial GPU memory and FLOPs, while SLM tuning distilled from LLMs can suffer from inefficient or low-quality supervision, leading to wasted LLM computation. SLM–LLM collaboration alleviates these by relocating supervision to low-cost venues and compressing guidance: (1) Selective Knowledge Distillation (SLM $\Leftarrow$ LLM, Fig. 6) distills selective LLM signals to reduce LLM calls; and (2) Proxy Transfer (SLM $\Rightarrow$ LLM, Fig. 7(a), Fig. 7(b)) generates lightweight, reusable updates for efficient LLM adaptation.
@@
Adaptive selection reduces data and teacher calls (LLKD Li et al. (2025a), efficiency-oriented KD Jazbec et al. (2024)); dense supervision uses concise yet rich signals such as contrastive label refinement (SynCID Liang et al. (2024)), compact CoT (HAWKEYE She et al. (2025)), and causal explanations Muhebwa and Osman (2025); targeted matching focuses on task-relevant features (C2KD Chen et al. (2025c)). Behavioral KD distills tool-use trajectories (Agent Distillation Kang et al. (2025)). Applications, e.g., LSC4Rec Lv et al. (2025), SmallPlan Pham et al. (2025), and Think Bergner et al. (2024), show that these strategies jointly cut FLOPs, tokens, and LLM calls while retaining near-LLM performance under tight compute and latency constraints.
@@
Models can train parameter updates on a small proxy model and lift them for use in the LLM. LoRAM Zhang et al. (2025a) trains low-rank adapters on an SLM to locate effective subspaces, later reused in LLMs. LiteMOE Zhuang et al. (2024) extracts proxy submodels for on-device tuning, enabling local specialization and feedback propagation. GateKeeper Rabanser et al. (2025) cascades an SLM and LLM to handle easy cases locally and defer harder ones, cutting inference cost, while logit-fusion Fan et al. (2025c) integrates fine-tuned SLM outputs to refine LLM distributions with minimal updates.
@@ 3.3 Collaboration During Inference Stage
Inference cost arises from three factors: (i) executing large models for every query (high FLOPs, latency, and API cost), (ii) decoding inefficiency (token-by-token verification by a single heavy model), and (iii) misallocated compute (overspending on easy queries, underspending on hard ones). SLM–LLM collaboration mitigates these through: (1) Cascade routing, handling easy queries with an SLM and escalating difficult ones (Fig. 8); (2) Speculative decoding, an SLM drafts tokens that a stronger verifier approves in bulk (Fig. 9); and (3) Compute-optimal test-time scaling, optimizing model selection and sampling budgets across SLM/LLM under fixed budget (Fig. 10).
@@
Figure 8: Cascade Routing
Query-level routers route per input using lightweight difficulty predictors Chen et al. (2024); Ong et al. (2025); Zhang et al. (2024d); Yue et al. (2024); Hao et al. (2024), e.g., SlimPM flags missing LLM knowledge for retrieval Tan et al. (2024). Token-level routers escalate only complex spans to LLMs, improving cost–quality trade-offs Zheng et al. (2024); Fu et al. (2025a). Resource-aware routers optimize selection via contextual bandits Wang et al. (2025c) or MoE-style edge–cloud load balancing Jin et al. (2025).
@@
Figure 9: Speculative Decoding
Speculative decoding advances focus on (i) decoupled draft–verify architectures, (ii) parallel decoding, and (iii) efficient deployment topologies. Cross-attention drafters (Beagle Zhong et al. (2025)) simplify verifier fusion; LookAhead Reasoning Fu et al. (2025b) and Adaptive Parallel Decoding Israel et al. (2025) improve parallelism; edge–cloud systems (LLMCAD Xu et al. (2023), PICE Zhan et al. (2025), FS-GEN Zhang et al. (2024e)) reduce verification cost via local drafting and adaptive policies.
@@
Figure 10: Compute-optimal TTS
@@ 3.4 Discussion
Existing cost-effectiveness collaborations show clear savings across compute, latency, communication, storage and API spend by coordinating the ML lifecycle, but remain constrained by fragmented pipelines, inconsistent cost reporting and brittle rules for when to hand off to the LLM Behera et al. (2025). Future work should standardize end-to-end cost-effectiveness metrics and use adaptive routing–paired with compression, bidirectional distillation etc, to deliver reliable savings across SLM-LLM collaboration paradigms.
@@ 4 Privacy-preserving Cloud–Edge Collaboration
Deploying LLMs on edge devices is limited by storage and computation constraints. A practical alternative is a cloud–edge architecture where lightweight SLMs handle local, context-aware tasks, and cloud-based LLMs provide broader generalization. However, transmitting sensitive data from trusted edge SLMs to semi-trusted cloud LLMs raises privacy concerns, as personally identifiable or proprietary information may be logged and leaked for training Duan et al. (2023), posing long-term security and compliance risks.
To balance privacy and performance, privacy-preserving collaboration has emerged at both the inference and fine-tuning stages. At the inference stage, SLMs act as Sensitive Information Gatekeepers, filtering or anonymizing data before sending it to LLMs, or as All-Information Guardians, retaining private context locally and integrating it after generation. At the fine-tuning stage, collaboration occurs through SLMs as Learners (learning from LLMs via distillation), LLMs as Learners (adapting via proxy tuning using SLM signals), or Collaborative Learners (jointly improving through federated knowledge transfer).
@@ 4.1 Inference Stage
During inference, edge user requests often contain sensitive information. Edge SLMs can process them locally but with lower quality than cloud LLMs, while direct cloud LLM use risks data exposure. Thus, privacy-preserving edge–cloud collaboration is needed, based on protection levels, categorized into (1) Sensitive Information Gatekeepers and (2) All-Information Guardians.
@@
(1) SLMs as Sensitive Information Gatekeepers To protect sensitive information in edge–cloud collaboration, the edge SLM serves as a privacy-preserving filter, for (i) sanitizing sensitive content and (ii) selectively disclosing necessary information before transmitting requests to the cloud.
(i) Sanitizing sensitive information: Fig.11(a) shows how SLMs remove or rephrase prompts containing personally identifiable information (PII). Chong et al. (2024) use an SLM to filter PII before cloud transmission, while Papadopoulou et al. (2022) employ SLM probabilities for detection. Uzor et al. (2025) rephrase queries to preserve meaning while removing PII, applied to clinical data Wiest et al. (2024). Hartmann et al. (2024) further enhance SLMs by masking sensitive data and using LLM-generated in-context examples.
(ii) Selective disclosure: Fig. 11(b) illustrates selective disclosure, where the edge SLM shares limited, task-relevant information to reduce privacy risks. Zhang et al. (2024f) share sketches or logits with the LLM, while CORE Fan et al. (2025a) ranks and filters relevant context. Zhang et al. (2024d) employ a policy network to balance privacy and quality, and Cheng et al. (2025) enforce a differential privacy budget to mathematically bound potential leakage. In sum, these methods shift from full-text disclosure to abstracted representation sharing, mitigating data exposure in SLM–LLM collaboration.
@@
Edge devices often enforce strict privacy rules that prevent PII disclosure. In such settings, SLMs serve as All-Information Guardians, handling sensitive data on-device while drawing on cloud LLM knowledge, ensuring privacy preservation (Fig. 12). Sketch-based CoGenesis Zhang et al. (2024f) illustrates this by using cloud LLMs for planning and local SLMs for personalized adaptation. Similar frameworks integrate cloud knowledge with local data for clinical Zhang et al. (2024g) or retrieval-augmented Cheng et al. (2025); Hartmann et al. (2024) applications. Industrial systems such as Apple Intelligence Gunter et al. (2024) follow the same principle, processing sensitive data locally and sharing only abstractions with the cloud.

### v2

@@ 3.1 Collaboration During Pre-Training Stage
The pre-training stage constitutes a computationally intensive phase of the LLM lifecycle, making it a primary target for cost-effectiveness optimization. The challenge lies in achieving high model capability without incurring the prohibitive expense of training on a massive yet low-quality data corpus. SLM–LLM collaboration during the pre-training stage addresses this by employing inter-model guidance to improve data efficiency and optimize the training trajectory. These methods can be categorized into three types: (1) Data curation, where Fig. 6(a) illustrates an LLM refining data for efficient SLM pre-training, and Fig. 6(b) shows an SLM evaluating data quality to reduce LLM pre-training cost; (2) Co-curriculum (Fig. 7(a)), which progressively grows from small to large model capacity; and (3) Pre-training distillation (Fig. 7(b)), which directly transfers knowledge from a teacher large model to a student small model.
(1) Data Curation
(a) SLM selects/reshapes data and feeds it into LLM
@@
High-quality data are vital for language model training, yet real-world corpora are often low-quality and noisy. Using low-quality data inflates computational costs and degrades model performance. SLM–LLM collaboration enhances data curation cost-effectively: (1) SLMs select or reshape data for LLM pretraining to boost performance with less data (Fig. 6(a)); (2) SLMs estimate data quality (e.g., perplexity) to guide LLM pretraining efficiently (Fig. 6(b)). For instance, the SALT Rawat et al. (2024) shows that a compact model can supply soft labels and identify informative hard examples, accelerating large-model pre-training while improving quality, reducing steps, FLOPs, and energy; Conversely, Ankner et al. (2024) adopts perplexity-based filtering with an SLM as the reference scorer to remove low-yield documents/tokens, cutting the optimization steps needed to reach a target loss.
@@
Figure 7: (a) Co-curriculum learning; (b) Knowledge distillation.
[…]
(3) Pre-Training Distillation
Pre-training distillation (PD) transfers knowledge from a teacher LLM to a student SLM through logit matching and data refinement, reducing data requirements while enhancing model capability (see Fig. 7(b)). Existing works fall into two categories based on the supervision signal: (i) logits-based and (ii) data-based pre-training distillation. For the former, task-agnostic distillation frameworks such as DistilBERT Sanh et al. (2019), TinyBERT Jiao et al. (2020), and MiniLM Wang et al. (2020) transmit teacher logits and embeddings to students as supervision signals. Peng et al. (2025) extend this to larger student LMs, addressing challenges in logits storage, compression, signal design, capacity gap, and online–offline trade-offs. DPT Goyal et al. (2025) further shows logits-based pre-training distillation enhances test-time scaling but weakens in-context learning, providing analytical insights into this trade-off. For the latter, MiniPLM Gu et al. (2025) performs offline teacher inference to reshape the pre-training corpus by (a) difficulty reweighting, prioritizing examples with informative teacher signals, and (b) diversity balancing, expanding topical, stylistic, and structural coverage, then trains a compact student on a refined mix.
@@ 3.2 Collaboration During Tuning Stage
Fine-tuning language models is costly along two fronts: (1) for tuning SLMs, relying on LLM distillation signals during adaptation increases FLOPs or API expenses, and fine-tuning quality depends on supervision quality: poor data wastes both tokens and computation and (2) directly fine-tuning the LLM is expensive in terms of FLOPs, memory, and time. SLM–LLM collaboration mitigates these costs by: (1) Selective Knowledge Distillation (SLM $\Leftarrow$ LLM) replaces repeated LLM usage with distilled SLMs and increases information per token via selective signals and sampling (Fig. 8); and (2) Proxy Transfer (SLM $\Rightarrow$ LLM) employs SLMs to produce reusable updates or control signals for efficient LLM adaptation (Figs. 9(a) and 9(b)).
@@
Full teacher supervision is costly and often redundant, while selective KD focuses limited guidance where it moves the needle most. We frame selective KD as teacher-in-the-loop curation that minimizes adaptation cost while preserving quality (Fig. 8). The pipeline selects high-yield samples where the teacher is confident but the student is uncertain, shapes supervision beyond hard labels to raise information per token, targets imitation to task-relevant representations instead of uniform mimicry, and then closes the loop using student feedback to guide the next round of teacher queries. In this loop, adaptive selection steers supervision to the error frontier—reducing data and teacher calls (LLKD Li et al. (2025a); student-in-the-loop selection with retained rationales Jazbec et al. (2024)); dense yet compact supervision replaces bare labels with short, information-rich signals that clarify boundaries and reasoning—contrastive label refinement (SynCID Liang et al. (2024)), concise CoT (HAWKEYE She et al. (2025), and causal explanations for robustness Muhebwa and Osman (2025); Lorè et al. (2024)); targeted matching aligns losses with where features live—layers/heads and cross-head logits that carry discriminative structure (C2KD Chen et al. (2025d)); and behavioral KD distills compact tool-use trajectories to amortize exploration (Agent Distillation Kang et al. (2025)). Applications instantiate the same loop end-to-end: LLM-assisted candidate generation with on-device ranking (LSC4Rec Lv et al. (2025)) operationalizes adaptive selection in recommendation systems; latency-sensitive recommenders pair KD with pruning/quantization Behdin et al. (2025) to shape supervision under strict budgets; lightweight planners distilled via LLM-guided SFT/RL (SmallPlan Pham et al. (2025)) emphasize targeted matching to task-relevant skills; and “single frozen-LLM prompt, SLM decode” amortization Bergner et al. (2024) closes the loop by throttling future teacher calls based on student performance—collectively cutting FLOPs, tokens, and LLM calls while retaining near-LLM quality.
@@
Proxy transfer reduces LLM adaptation costs by using small models as proxies to generate parameters or signals that can be reused by the larger model. This strategy operates through two main pathways: (1) SLMs, often compressed from the LLM, are fine-tuned with low resources to produce LoRA-like updates compatible with the original LLM (Fig. 9(a)); and (2) fine-tuned SLMs have their outputs merged with those of the LLM, simulating the effect of fine-tuning the large model through a lightweight proxy (Fig. 9(b)). Both pathways reduce the LLM’s required FLOPs, data, and tuning time. For example, LoRAM Zhang et al. (2025a) trains low-rank adapters on an SLM to locate effective subspaces, later reused in LLMs. LiteMOE Zhuang et al. (2024) extracts proxy submodels for on-device tuning, enabling local specialization and feedback propagation. GateKeeper Rabanser et al. (2025) cascades an SLM and LLM to handle easy cases locally and defer harder ones, cutting inference cost, while G-Boost Fan et al. (2025c) integrates fine-tuned SLM outputs to refine LLM distributions with minimal updates.
@@ 3.3 Collaboration During Inference Stage
Inference cost arises from three sources: (i) running large models on every request (high FLOPs, latency, or API cost); (ii) decoding inefficiency (token-by-token verification by a single heavy model); and (iii) misallocated compute (overspending on easy queries and underspending on hard ones). SLM–LLM collaboration at the inference stage can address these through: (1) Cascade routing: a lightweight router which is guided using some uncertainty threshold, sending most traffic to a cheap SLM and escalating only when necessary (Fig. 10); (2) Speculative decoding: a fast drafter (often an SLM) proposes multiple tokens that a strong verifier (LLM) accepts or corrects in bulk (Fig. 11); and (3) Compute-optimal test-time scaling (TTS): under a fixed budget, optimizes model choice and sampling budgets across SLM/LLM families to meet the target quality (Fig. 12).
@@
Figure 10: Cascade Routing.
Cascade routing reduces average serving costs by directing most requests to a cheap SLM and escalating only complex cases to an LLM. Routers operate at different granularities: (i) query-level for holistic request classification, (ii) token/span-level for fine-grained generation control, and (iii) resource-aware for expert selection under system constraints.
(i) Query-level routers make a single deferral decision per input. Classical methods train a lightweight classifier to judge response reliability Chen et al. (2024a); Ong et al. (2025); Zhang et al. (2024d). Extensions predict query difficulty for adaptive routing Yue et al. (2024); Hao et al. (2024), including SlimPM Tan et al. (2024), which uses a slim proxy to detect an LLM’s knowledge gaps and decide when to retrieve.
(ii) Token/span-level routers exert finer control by only invoking the LLM for critical tokens or segments. This improves the cost–quality trade-off by letting SLMs handle non-critical generation and LLMs only for critical tokens (e.g., CITER Zheng et al. (2024) and Roads to Rome Fu et al. (2025b)).
(iii) Resource-aware routers frame model selection as a constrained optimization problem, which uses contextual bandits to pick the best model under latency constraints Wang et al. (2025c) or use MoE gates to dispatch requests across edge/cloud models, balancing load and budget Jin et al. (2025).
@@
Figure 11: Speculative Decoding.
Speculative decoding reduces LLM compute by allowing a cheaper drafter to propose tokens (or short continuations) that a stronger verifier accepts in bulk (Fig. 11). Existing studies can be grouped into three major directions: (i) adaptive draft–verify pipelines, (ii) system- and hardware-aware parallelization, and (iii) distributed and hierarchical multi-model collaboration.
(i) Adaptive Draft–Verify Pipelines Early frameworks such as speculative decoding Leviathan et al. (2023); Chen et al. (2023) and SpecDec Xia et al. (2023) establish the draft–then–verify paradigm, where SLMs propose multi-token drafts and LLMs verify them in parallel for speedups. BiLD Kim et al. (2023), staged SD Spector and Re (2023), and LLMCad Xu et al. (2023) extend this idea to big-little or on-device settings with rollback and correction strategies. Later methods including SpecInfer Miao et al. (2024), cascade SD Chen et al. (2024b), and SpecTr Sun et al. (2023), introduce tree- or cascade-based verification, while online SD Liu et al. (2024c), TurboSpec Liu et al. (2024b), SCD Yuan et al. (2024), and mixture-of-attentions SD Zimmer et al. (2024) add adaptive retraining, feedback control, and enhanced drafting accuracy. Together, these approaches evolve SD into dynamic, feedback-driven pipelines that balance quality and efficiency.
(ii) System- and Hardware-Aware Parallelization Several systems accelerate SD through pipeline and hardware optimization. EdgeLLM Xu et al. (2024b), SpecExec Svirschevski et al. (2024), and SPIN Chen et al. (2025b) boost throughput via multi-token speculation, GPU pipelining, and heterogeneous SLM selection, while Dovetail Zhang et al. (2024g) and APD Israel et al. (2025) exploit heterogeneous devices and adaptive speculative width. Hardware-level efforts like HADES Yang et al. (2025) embed SD primitives into accelerators, and HiSpec Kumar et al. (2025) uses early-exit layers and cache reuse for lightweight intermediate verification. These methods make SD a pipeline-optimized, hardware-co-designed inference paradigm.
(iii) Distributed and Hierarchical Multi-Model Collaboration Edge–cloud and multi-agent extensions broaden SD into large-scale collaboration. Fast edge–cloud SD Venkatesha et al. (2025), SLED Li et al. (2025e), and heterogeneous SD frameworks Zhu et al. (2025) distribute drafting to edge SLMs and verification to cloud LLMs using asynchronous batching. SpecServe Huang et al. (2025b) and CoSine Gao et al. (2025a) add adaptive schedulers and confidence-prior verifiers for SLO-aware efficiency, while PICE Zhan et al. (2025) and CoS Fu et al. (2025a) extend SD to semantic-level and multi-model collaboration through dynamic task routing and alternating proposer–verifier roles. Together, these systems turn SD into a distributed orchestration framework for scalable, low-cost LLM inference.
@@
Figure 12: Compute-optimal TTS.
@@ 3.4 Discussion
Existing cost-effectiveness collaborations demonstrate clear savings across compute, latency, communication, storage, and API costs by coordinating the ML lifecycle but remain limited by fragmented pipelines, inconsistent cost reporting, and brittle rules for determining when to hand off to the LLM Behera et al. (2025). Future work should standardize end-to-end cost-effectiveness metrics and employ adaptive routing, paired with compression, bidirectional distillation, and related techniques, to ensure reliable savings across SLM–LLM collaboration paradigms.
@@ 4 Privacy-preserving Cloud–Edge Collaboration
Deploying LLMs on edge devices is limited by hardware constraints. A practical alternative is a cloud–edge framework where lightweight SLMs handle local, context-aware tasks, and cloud-based LLMs provide broader generalization Lv et al. (2023); Shao et al. (2025). However, transmitting sensitive data from trusted edge SLMs to semi-trusted cloud LLMs raises privacy concerns, as personally identifiable or proprietary information may be logged and leaked for training Duan et al. (2023), posing long-term data security risks.
To balance privacy and performance, privacy-preserving cloud-edge collaboration has emerged at both the inference and fine-tuning stages. At the inference stage, SLMs act as (1) Sensitive Information Gatekeepers, filtering or anonymizing data before sending it to LLMs, or as (2) All-Information Guardians, retaining private context locally and integrating it after generation. At the fine-tuning stage, collaboration occurs via (1) SLMs as Learners (learning from LLMs via distillation), (2) LLMs as Learners (adapting via proxy tuning using SLM signals), or (3) Collaborative Learners (jointly improving through federated knowledge transfer).
@@ 4.1 Inference Stage
During inference, edge user requests may contain sensitive information. Edge SLMs can process locally but with lower quality than cloud LLMs, while direct cloud LLM use risks data exposure. Thus, privacy-preserving edge–cloud collaboration is needed, based on protection levels, categorized into (1) Sensitive Information Gatekeepers (Fig. 13) and (2) All-Information Guardians (Fig. 14).
@@
(1) SLMs as Sensitive Information Gatekeepers
To protect sensitive information in edge–cloud collaboration, the edge SLM serves as a privacy-preserving filter, for (i) sanitizing sensitive content and (ii) selectively disclosing necessary information before transmitting requests to the cloud.
(i) Sanitizing sensitive information: Fig. 13(a) shows how SLMs remove or rephrase prompts containing personally identifiable information (PII). Casper Chong et al. (2024) use an SLM to filter PII before cloud transmission, while Papadopoulou et al. (2022) employ SLM probabilities for detection. LLM gatekeeper Uzor et al. (2025) rephrases queries to preserve meaning while removing PII, applied to clinical data Wiest et al. (2024). Hartmann et al. (2024) enhance SLMs by masking sensitive data and using LLM-generated examples.
(ii) Selective disclosure: Fig. 13(b) illustrates selective disclosure, where the edge SLM shares limited, task-relevant information to reduce privacy risks. CoGenesis Zhang et al. (2024f) shares sketches or logits with the LLM, while CORE Fan et al. (2025a) ranks and filters relevant context. P3Defer Zhang et al. (2024d) employs a policy network to balance privacy and quality, and RemoteRAG Cheng et al. (2025) enforces a differential privacy budget to mathematically bound potential leakage. In sum, these methods shift from full-text disclosure to abstracted representation sharing, mitigating data exposure in SLM–LLM collaboration.
@@
Edge device policies may impose strict privacy requirements that prohibit any PII disclosure. Under such constraints, edge SLMs act as All-Information Guardians, processing sensitive data entirely on-device while integrating cloud LLM knowledge, ensuring that private information never leaves the user’s control, as shown in Fig. 14. Sketch-based CoGenesis Zhang et al. (2024f) exemplifies the paradigm, using the cloud LLM for high-level planning while the SLM personalizes results locally with user logs. PrivacyBoost-SLM Zhang et al. (2024i) applies this to clinical decision-making, combining cloud-derived medical knowledge with patient data. RemoteRAG Cheng et al. (2025) extends this idea by pairing SLMs with local retrievers (RAG) to integrate private and domain-general knowledge without exposing raw data. Hartmann et al. (2024) further employ LLMs as RAG example providers to enhance the SLM context. Industrial systems such as Apple Intelligence Gunter et al. (2024) adopt a similar approach, processing sensitive data on-device and sending only abstracted information to the cloud.